### v1

# Historically Relevant Event Structuring for Temporal Knowledge Graph Reasoning

Abstract.
Temporal Knowledge Graph (TKG) reasoning focuses on predicting events through historical information within snapshots distributed on a timeline. Existing studies mainly concentrate on two perspectives of leveraging the history of TKGs, including capturing evolution of each recent snapshot or correlations among global historical facts. Despite the achieved significant accomplishments, these models still fall short of (1) investigating the influences of multi-granularity interactions across recent snapshots and (2) harnessing the expressive semantics of significant links accorded with queries throughout the entire history, especially events exerting a profound impact on the future. These inadequacies restrict representation ability to reflect historical dependencies and future trends thoroughly. To overcome these drawbacks, we propose an innovative TKG reasoning approach towards Historically Relevant Events Structuring ($\mathsf{HisRES}$). Concretely, $\mathsf{HisRES}$ comprises two distinctive modules excelling in structuring historically relevant events within TKGs, including a multi-granularity evolutionary encoder that captures structural and temporal dependencies of the most recent snapshots, and a global relevance encoder that concentrates on crucial correlations among events relevant to queries from the entire history. Furthermore, $\mathsf{HisRES}$ incorporates a self-gating mechanism for adaptively merging multi-granularity recent and historically relevant structuring representations. Extensive experiments on four event-based benchmarks demonstrate the state-of-the-art performance of $\mathsf{HisRES}$ and indicate the superiority and effectiveness of structuring historical relevance for TKG reasoning.
Temporal Knowledge Graph, Graph Neural Network, Extrapolation, Link Prediction
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Temporal reasoning
1. Introduction
Temporal Knowledge Graphs (TKGs) (Cai et al., 2022) have revolutionized the limitations of static KGs by considering temporal information. Specifically, TKGs advance beyond traditional triple facts, enhancing them into (subject, relation, object, timestamp) quadruples, thereby bridging the divide between static KGs and dynamic characteristics. This enrichment facilitates capturing the dynamic evolution of real-world events and holds significant practical relevance. However, TKGs still suffer from incompleteness and encounter unique issues due to their distinct temporal nature. This pressing task in TKGs can be divided into interpolation and extrapolation, involving reasoning missing facts (events) at past and future timestamps according to historical knowledge, respectively. Notably, this paper primarily focuses on the extrapolation task that has been widely explored in recent years. It expands the downstream application domains of KGs to include event prediction (Zhou et al., 2023), social network analysis (Cai et al., 2018), sensor fault detection (Cai and Zheng, 2020), etc.
As outlined in prior research (Han et al., 2021; Li et al., 2021b), TKGs manifest in two primary forms. One form is a multi-relational graph akin to static KGs but enhanced with timestamped relation edges to convey dynamic characteristics, while the other can be represented as sequences of snapshots or subgraphs partitioned by timeline. In the latter form, each snapshot encapsulates concurrent events transpiring at specific timestamps, offering a more intuitive depiction of the evolutionary patterns of events.
In the extrapolation task, several works introduce novel perspectives that emphasize different aspects of historical information within TKGs. According to the views and ideas, they can be divided into two categories: (a) tallying statistics of historical events and leveraging repetitive patterns to predict future facts (Zhu et al., 2021; Xu et al., 2023); (b) capturing concurrent and evolutionary correlations from the most recent snapshots (Jin et al., 2020; Li et al., 2021b, 2022a; Liu et al., 2023; Liang et al., 2023; Zhang et al., 2023a). In detail, the (a) category utilizes simple linear transformations as encoders and generates sparse matrices from repetitive statistics to mask predictions, while the (b) category adopts Graph Neural Networks (GNNs) to model the structural dependencies among concurrent facts, and recurrent units such as Gated Recurrent Unit (GRU) to capture the sequential evolution of snapshots.
Based on the (b) category, few innovative models (Li et al., 2022c; Zhang et al., 2023b; Chen et al., 2024) additionally incorporate information from the entire history to achieve more comprehensive learning of TKGs. Within them, TiRGN (Li et al., 2022c) incorporates a global historical encoder like CyGNet (Zhu et al., 2021) to constrain the prediction scope, HGLS (Zhang et al., 2023b) establishes connections between common entities across different timestamps to capture multi-hop correlations among globally historical events, LogCL (Chen et al., 2024) constructs a graph consisting of historical facts relevant to queries to obtain rich semantic information. However, despite these advancements, the above studies still fall short of comprehensively exploring historically relevant events in TKGs. It can be analyzed into two specific challenges: (1) investigating the influences of multi-granularity interactions across recent snapshots and (2) harnessing the expressive semantics of significant links accorded with queries throughout the entire history, especially events exerting a profound impact on the future.
Figure 1. An illustration of extrapolation over TKGs.
Specifically, the first noteworthy challenge arises due to the lack of explicit and flexible judgment on multi-granularity correlations between facts or entities across adjacent snapshots. Since many sequentially occurring real-world events at adjacent timestamps are causally linked, the above-mentioned deficiency leads to the loss of crucial dependencies between such events from adjacent timestamps. It could originally provide valuable clues for TKG reasoning. While RPC (Liang et al., 2023) and LogCL (Chen et al., 2024) involve inter-snapshot information and assign different weights to distinguish contributions of recent snapshots, they remain limited to the snapshot level and may overlook specific links within snapshots. A vivid example from ICEWS (Boschee et al., 2015) dataset is illustrated in Figure 1, (North_America, Host_a_visit, Business_(Africa)) is caused by (Barack_Obama, Consult, North_America) at the previous time. Instead of complex evolutionary patterns, we can capture such correlations with simple 2-hop links (denoted in red) through structuring and learning events inter-snapshot.
For the second significant challenge, prior approaches tend to neglect valuable information intrinsic in the structure of globally relevant events, such as long-term semantics and periodic interactions, leading to inadequate representations. In Figure 1, it is evident that the triple ($e_{o}$, $r_{0}$, $e_{1}$) in $\mathcal{G}_{0}$ holds profound implications for future events. By identifying significant facts relevant to queries like this and structuring links (denoted in blue) throughout the entire history, we can effectively leverage crucial events from a global perspective. Although HGLS (Zhang et al., 2023b) proposes to connect common entity in each snapshot through additional edges to capture entity dependencies, it may incorporate redundant information from distant timestamps while necessitating more hops and layers for modeling neighboring interactions. LogCL (Chen et al., 2024) provides a similar idea for integrating historical facts relevant to queries, but it lacks the means to prioritize key historical links, leading to reduced efficiency.
To tackle these aforementioned challenges, this paper proposes a novel method that specializes in TKG reasoning with Historically Relevanct Event Structuring ($\mathsf{HisRES}$). Specifically, $\mathsf{HisRES}$ adopts an encoder-decoder architecture, focusing on two critical aspects of TKG reasoning: capturing the structural and temporal patterns within recent snapshots and understanding the impact of relevant events throughout the entire history. In terms of recent history, $\mathsf{HisRES}$ employs a multi-granularity evolutionary encoder (§3.2), which models concurrent interactions and evolution within each snapshot and further integrates adjacent snapshots to learn temporal dependencies using multi-hop links (addressing the first challenge). To explicitly capture structural correlations across the entire history (addressing the second challenge), $\mathsf{HisRES}$ introduces a global relevance encoder, structuring historical facts directly relevant to the query set. Specially, we propose an innovative convolution-based graph attention network (ConvGAT) to efficiently emphasize important links of historical relevance for future prediction. Furthermore, $\mathsf{HisRES}$ incorporates a self-gating mechanism (§3.3) to adaptively fuse entity representations from different granularities or encoders, achieving a comprehensive understanding by considering various historical information (i.e., global and local, intra-snapshot and inter-snapshot) of TKGs, leading to accurate predictions. Extensive experiments on four benchmarks demonstrate the superiority and effectiveness of $\mathsf{HisRES}$ in TKG reasoning, particularly in structuring historical relevance.
In general, the main contributions of this paper can be summarized as follows:
  • We propose a novel TKG reasoning method termed $\mathsf{HisRES}$, which sufficiently discovers structural and temporal dependencies within TKGs through historically relevant event structuring.
  • Through the adaptive fusion of multi-granularity evolutionary and global relevance encoding, $\mathsf{HisRES}$ can comprehensively integrate representations of TKGs throughout historical correlations.
  • Extensive experiments on four event-based TKG datasets demonstrate that $\mathsf{HisRES}$ outperforms state-of-the-art TKG reasoning baselines, showcasing its effectiveness and superiority.
2. Related Works
KG reasoning (Tian et al., 2021) has been extensively studied in recent years, making significant contributions to downstream applications. Traditional reasoning deals mainly with static KGs such as the widely used FB15k-237 (Toutanova et al., 2015) and WN18RR (Dettmers et al., 2018). In contrast, TKG reasoning is tailored for time-sensitive data, adept at capturing evolutionary patterns, and discovering future facts within event-based datasets characterized by temporal features.
Figure 2. An illustration of $\mathsf{HisRES}$ model architecture.
2.1. Static Knowledge Graph Reasoning
Early research on static KGs focuses on modeling entities and relations in various representation spaces. TransE (Bordes et al., 2013) introduces a simple translation assumption $s+r\approx o$, while DistMult (Yang et al., 2015) learns relational semantics through embeddings from a bilinear objective. In the complex vector space, ComplEx (Trouillon et al., 2016) and RotatE (Sun et al., 2019) employ Hermitian dot products and relation rotations, respectively. With the advent of convolutional neural networks, approaches such as ConvE (Dettmers et al., 2018) and ConvTransE (Shang et al., 2019) utilize convolutional operations to fuse entity-relation embeddings. Recent advancements aim at learning more expressive structures with Graph Convolutional Networks (GCN) (Kipf and Welling, 2017). R-GCN (Schlichtkrull et al., 2018) extending GCNs with relation-specific block-diagonal matrices. CompGCN (Vashishth et al., 2020) integrates multi-relational GCN methods with various composition operations, while KBGAT (Nathani et al., 2019) introduces an attention-based mechanism to capture neighboring semantics. However, these static approaches fail to incorporate temporal information effectively, making it hard to handle dynamic facts or learn evolutionary patterns.
2.2. Temporal Knowledge Graph Reasoning
According to the occurrence moment of predicted events (i.e., known or future timestamps), TKG reasoning can be roughly divided into two types: interpolation and extrapolation. Notably, this paper focuses on the extrapolation task.
Assuming we know all facts before $t_{T}$ (including $t_{T}$), the extrapolation task aims to predict future facts at timestamps $t>t_{T}$ based on historical TKG sequences. Various approaches have been explored in recent years to address it. xERTE (Han et al., 2021) designs a subgraph sampling module based on temporal information and incorporates time within relation embeddings. TITer (Sun et al., 2021) and CluSTER (Li et al., 2021a) employ reinforcement learning to generate query-specific paths composed of historical entities. Besides, numerous studies can be categorized into two main groups: those that focus on repeating statistical historical facts, and those that model the interactions among historical events. In the first category, CyGNet (Zhu et al., 2021) utilizes multi-hot vector as a historical vocabulary that records facts from the past and masks prediction. Building upon this, CENET (Xu et al., 2023) introduces contrastive learning and learnable classifiers to differentiate historical events. In the latter category, RE-NET (Jin et al., 2020) and RE-GCN (Li et al., 2021b) employ GCNs to model concurrent structures and utilize recurrent units to learn entity or relation evolution patterns. On this basis, CEN (Li et al., 2022a) considers historical length-diversity and time-variability. TiRGN (Li et al., 2022c) adopts a global history encoder like CyGNet to identify entities related to historical or unseen facts and a time-guided decoder. RETIA (Liu et al., 2023) and RPC (Liang et al., 2023) introduce the concept of line graphs to learn the interaction of relations. L2TKG (Zhang et al., 2023a) addresses the missing associations and creates links between unconnected entities to obtain more comprehensive representations. Some novel insights attempt to capture global structural correlations through additional graphs. HGLS (Zhang et al., 2023b) explicitly captures long-term dependencies by connecting the same entity at each timestamp, while LogCL (Chen et al., 2024) proposes an entity-aware attention mechanism to flexibly adopt facts relevant to queries and utilizes contrastive learning to improve robustness. Nevertheless, none of these methods consider the multi-granularity interactions among the most recent snapshots, which are crucial for predicting future facts. Nor do they efficiently learn correlations from the entire history, resulting in somewhat excessive consumption, such as increasing the scale of the graph or introducing unnecessary facts.
3. Method
The overarching framework of $\mathsf{HisRES}$ is depicted in Figure 2, which follows encoder-decoder architecture and consists of three key components: (i) Multi-granularity Evolutionary Encoder3.2), which focuses on capturing correlations among concurrent facts in the most recent snapshots and modeling temporal evolution of these snapshots; (ii) Self-gating Mechanism3.3), which studies a learnable weight for each entity and adaptively merges entity representations from different encoders; (iii) Global Relevance Encoder3.4), which integrates all historically relevant facts, and aggregates their crucial neighboring semantics with a novel graph attention network.
3.1. Notations
In this paper, Temporal Knowledge Graphs (TKGs) are conceptualized as sequences of subgraphs, referred to as snapshots, and denoted as $G=\{\mathcal{G}_{1},\mathcal{G}_{2},...,\mathcal{G}_{\mathcal{T}}\}$. Here, $\mathcal{T}$ represents the total number of historical timestamps, and each $\mathcal{G}_{t}$ corresponds to the collection of facts or events occurring at timestamp $t$. All historical facts correlated with query pairs can be grouped within a global graph $\mathcal{G}^{\mathcal{H}}_{t}$ in $\mathsf{HisRES}$ when making predictions, such as $(s,r,?)$ at timestamp $t$. Given $\mathcal{G}=\{\mathcal{E},\mathcal{R},\mathcal{F},\mathcal{T}\}$, where $\mathcal{E}$, $\mathcal{R}$, and $\mathcal{F}$ represent the sets of entities, relations, and facts, respectively. Each fact $f\in\mathcal{F}$ is defined as a quadruple $(s,r,o,t)$, where $s$ and $o$ belong to $\mathcal{E}$, representing the subject and object entities, $r\in\mathcal{R}$ represents the relation, and $t\in\mathcal{T}$ indicates the timestamp of the quadruple. In this paper, bold letters are used to indicate vectors. In summary, the important notations is presented in Table 1.
Table 1. The Summary of important notations.
Notations Descriptions
$\mathcal{E}$, $\mathcal{R}$, $\mathcal{T},\mathcal{Q}$ entity, relation, timestamp, and query set of TKGs
$\mathcal{G}_{t}$ temporal knowledge graph snapshot at $t$
$\textbf{E}_{t}$, $\textbf{R}_{t}$ entity and relation embeddings of local evolution at $t$
$\mathcal{G}^{\mathcal{H}}_{t}$ globally relevant graph containing all facts before $t$
$\textbf{E}^{\mathcal{H}}_{t}$, $\textbf{R}^{\mathcal{H}}_{t}$ entity and relation embeddings of global relevance at $t$
[…]
$\Theta$ self-gating vectors for embedding fusion
$\alpha$ trade-off hyper-parameter for prediction task coefficient
$l$ hyper-parameter to control local historical length
Task Definition. Our primary focus is on the extrapolation task within TKG reasoning, specifically, entity prediction in future timestamps. Given a query $q=(s,r,?,t)$ or $q=(?,r,o,t)$, where $q\in\mathcal{Q}$, our learning objective is to calculate the conditional probability of an object $o$ or subject $s$ given $(s,r,t)$ or $(r,o,t)$ and past snapshots. This is denoted as $p(o|s,r,t,\mathcal{G}_{0:t-1})$ or $p(s|r,o,t,\mathcal{G}_{0:t-1})$. It is noteworthy that we utilize object prediction as an example in our methodology section.
3.2. Multi-granularity Evolutionary Encoder
Multi-granularity evolutionary encoder serves as the cornerstone of $\mathsf{HisRES}$, which entails predicting future events using historical snapshots from the most recent past. In accordance with this principle, we consider the most neighboring $l$ snapshots as historical data and aggregate features from both structural and temporal perspectives in the multi-granularity level. It includes intra-snapshot and inter-snapshot dependencies, the latter of which can structure several adjacent snapshots simultaneously (we select every two adjacent snapshots). The illustration of one cycle in multi-granularity evolutionary encoder is depicted in Figure 3, RU denotes relation updating.
Figure 3. The achitecture of proposed multi-granularity evolutionary encoder.
3.2.1. Spatial-temporal Modeling intra-snapshot
The fundamental component of $\mathsf{HisRES}$ involves modeling spatial and temporal features from each recent snapshot, encompassing all characteristics of TKGs (i.e., entity, relation and time information), which is shown in the intra-snapshot of Figure 3.
Time Encoding
Analyzing the evolution process of events, periodicity and time attributes are crucial factors. Thus, similar to previous works (Li et al., 2022b; Chen et al., 2024), we encode the time information $\Delta\textbf{t}$ based on the time interval between the current time $t_{i}$ and the predicted time $t$, as follow:
(1) $\Delta\textbf{t}=\sigma(\textbf{w}_{t}(t-t_{i})+\textbf{b}_{t}){,}$
where $\sigma$ is the cosine periodic activation function, $\textbf{w}_{t}$ and $\textbf{b}_{t}$ are the learnable parameter and bias of time. Hence, the initial entity embeddings for each single snapshot can be calculated as:
(2) $\textbf{E}^{g^{\prime}}_{t-1}=\textbf{W}_{0}([\textbf{E}^{g}_{t-1}||\Delta\textbf{t}]){,}$
where $\textbf{W}_{0}\in\mathbb{R}^{d\times 2d}$ is a linear transformation. In particular, all our modules use $t-1$ as a reference time.
Entity Aggregation
In the spatial aspect, the basic aggregation process resembles that of CompGCN (Vashishth et al., 2020). To merge entities and relations effectively, we employ composition operators such as ”subject+relation”, akin to RE-GCN (Li et al., 2021b). The formulation is as follows:
(3) $\textbf{o}^{(l+1)}_{t-1}=\sigma(\sum_{(s,r,o)\in\mathcal{F}_{t-1}}\textbf{W}^{(l)}_{1}(\textbf{s}^{(l)}_{t-1}+\textbf{r}^{(l)}_{t-1})+\textbf{W}^{(l)}_{2}\textbf{o}^{(l)}_{t-1})\textnormal{,}$
where $\mathcal{F}_{t-1}$ stands for the set of facts within the $t-1$ snapshot, $\textbf{s}_{t-1},\textbf{o}_{t-1}\in\textbf{E}^{g^{\prime}}_{t-1}$, $\textbf{r}_{t-1}\in\textbf{R}_{t-1}$, $\mathbf{W}^{(l)}_{1}$ and $\mathbf{W}^{(l)}_{2}$ are linear transformations for $(s,r)$ and the self-loop of objects in this layer, $\sigma$ represents the Randomized leaky Rectified Linear Unit (RReLU) activation function. The output of aggregated entity representation intra-snapshot is denoted as $\vec{\textbf{E}}_{t-1}$.
Entity Evolution
In the temporal aspect, we incorporate the GRU cell to progressively capture evolving entity representations from each recent snapshot to fit future facts, which can be expressed as:
(4) $\textbf{E}^{g}_{t}=\mathrm{GRU}(\vec{\textbf{E}}_{t-1},\textbf{E}^{g^{\prime}}_{t-1})\textnormal{.}$
Relation Updating
Additionally, $\mathsf{HisRES}$ enriches the evolution of relations and simultaneously updates the relation representation in CompGCN during entity aggregation. The operation is as follows:
(5) $\textbf{R}^{(l+1)}_{t-1}=\sigma(\textbf{W}^{(l)}_{r}\textbf{R}^{(l)}_{t-1})\textnormal{,}$
where $\sigma$ is RReLU activation function and $\textbf{W}^{(l)}_{r}$ denotes a linear layer. $\vec{\textbf{R}}_{t-1}$ represents the output relation representation of CompGCN. Finally, the relation embeddings $\textbf{R}_{t}$ are updated similarly to entities:
(6) $\textbf{R}_{t}=\mathrm{GRU}(\vec{\textbf{R}}_{t-1},pooling(\textbf{E}^{\mathcal{R}}_{t-1}))\textnormal{,}$
where pooling denotes the mean pooling operation, $\textbf{E}^{\mathcal{R}}_{t-1}$ denotes the embeddings of $\mathcal{R}$ related entities at $t-1$ snapshot.
3.2.2. Spatial-temporal Modeling inter-snapshot
In addition to modeling spatial-temporal correlations individually for each nearby snapshot, there are a large number of facts in TKGs that have a sequential correlation. As an example in the ICEWS dataset, (Citizen (Malaysia), Make optimistic comment, Ministry (Malaysia), t) is actually a follow-up event to (Ministry (Malaysia), Responses, Citizen (Malaysia), t-1). In this instance, the prior module designed for each individual snapshot fails to capture correlations such as this from a simple two-hop link, causing a complex mode.
To tackle this issue, we merge every two consecutive snapshots into a unified graph, which enables us to capture multi-hop association inter-snapshot using the same aggregation function as the former module (without sharing parameters). The diagram can be seen in the inter-snapshot of Figure 3. We omit the aggregation step, and denote the output of this process as $\textbf{E}^{gg^{\prime}}_{t-2:t-1}$. The evolutionary modeling proceeds as follows:
(7) $\textbf{E}^{gg}_{t}=\mathrm{GRU}(\textbf{E}^{gg^{\prime}}_{t-2:t-1},\textbf{E}^{gg}_{t-2:t-1})\textnormal{.}$
Since the graphs in this module are composed of adjacent snapshots, we do not employ relation updating as described in equation (5) of CompGCN, nor the time encoding as described in equation (2).
3.3. Self-gating Mechanism
$\mathsf{HisRes}$ incorporates a self-gating mechanism to adaptively merge multi-granularity entity embeddings from the aforementioned encoder. Given that we capture structure and evolution patterns across snapshots at various granularities, it is crucial for $\mathsf{HisRes}$ to effectively combine these features to comprehend correlations among recent snapshots. The self-gating is denoted as $\Theta\in\mathbb{R}^{|\mathcal{E}|\times d}$, which regulates a learnable weighting of different recent entity representations. Formally,
(8) $\textbf{E}_{t}=\Theta^{\mathcal{E}}\textbf{E}^{g}_{t}+(1-\Theta^{\mathcal{E}})\textbf{E}^{gg}_{t}\textnormal{,}$
(9) $\Theta^{\mathcal{E}}=\sigma(\textbf{W}_{3}\textbf{E}^{g}_{t}+\textbf{b}_{g}){,}$
where $\sigma(\cdot)$ is the Sigmoid activation function to limit the weight of $\textbf{E}^{g}_{t}$ and $\textbf{E}^{gg}_{t}$ between $[0,1]$, $\textbf{W}_{3}\in\mathbb{R}^{d\times d}$ is a linear transformation layer and $\textbf{b}_{g}$ denotes bias.
3.4. Global Relevance Encoder
While the former modules adeptly capture the most recent evolutionary patterns from a multi-granularity perspective, it falls short in highlighting pivotal events closely related to queries within the entire historical TKGs. These events can significantly impact future trends and deserve special consideration. Unlike the widely used method of making statistics from historically related facts (Zhu et al., 2021; Li et al., 2022c; Xu et al., 2023), we incorporate an effective approach to establish a globally relevant graph that amalgamates historically relevant events. This enables explicit modeling of correlations among crucial entities and leverages comprehensive historical interactions. Furthermore, it is worth noting that our strategy, compared to recent methods (Zhang et al., 2023b; Chen et al., 2024) that construct similar global graphs to utilize historical information, achieves an even more effective capture of global correlations while requiring fewer nodes and connections.
3.4.1. Global Relevance Structuring
The idea of structuring global relevance can be seen as an expansion of historical statistics. Our method involves extracting all related facts $f_{q}\in{\mathcal{G}_{0:t-1}}$ with query pairs $(s,r)$ from the entire historical TKGs, where $q\in\mathcal{Q}_{t}$. Subsequently, the globally relevant graph $\mathcal{G}^{\mathcal{H}}_{t}$ is constructed based on these query-specific facts. This structuring process is illustrated in Figure 4.
Figure 4. A construction illustration of the globally relevant graph. In the historical TKGs, query-specific facts are framed by green ovals, while the globally relevant graph comprises all those facts.
3.4.2. ConvGAT
We consider using the fusion of multi-granularity evolutionary encoder from self-gating as the initial entity embeddings in global relevance aggregation. Since our objective of global relevance encoder is to figure out the important correlations from the whole historically relevant facts, we design a convolution-based graph attention network called ConvGAT, which integrates a 1-D convolutional operator to merge entities and relations, and an attention mechanism to focus on those crucial correlations. Firstly, we calculate the attention scores $\alpha_{o,s}$ of each edge in $\mathcal{G}^{\mathcal{H}}_{t}$, which as follows:
(10) $\theta^{(l)}_{o,s}=\frac{exp(\textbf{W}^{(l)}_{4}\sigma(\textbf{W}^{(l)}_{5}[\textbf{s}^{(l)}||\textbf{r}||\textbf{o}^{(l)}]))}{\sum_{s^{\prime}\in\mathcal{N}_{(o)}}exp(\textbf{W}^{(l)}_{4}\sigma(\textbf{W}^{(l)}_{5}[\textbf{s'}^{(l)}||\textbf{r}||\textbf{o}^{(l)}])}\textnormal{,}$
where $\textbf{s},\textbf{o}\in\textbf{E}_{t}$, $\textbf{W}^{(l)}_{4}\in\mathbb{R}^{3d}$ and $\textbf{W}^{(l)}_{5}\in\mathbb{R}^{3d\times 3d}$ represent linear transformation layers, $\sigma$ corresponds to the LeakyReLU activation function, $\mathcal{N}_{(o)}$ denotes the neighboring subjects of object $o$. Notably, we do not update relation embeddings in global relevance encoder. Subsequently, the aggregation function of global relevance encoder is defined as follows:
(11) $\textbf{o}^{(l+1)}=\sigma(\sum_{(s,r,o)\in\mathcal{G}^{\mathcal{H}}_{t}}\theta^{(l)}_{o,s}\textbf{W}^{(l)}_{6}\psi(\textbf{s}^{(l)}+\textbf{r})+\textbf{W}^{(l)}_{7}\textbf{o}^{(l)})\textnormal{,}$
where $\sigma$ is the RReLU activation function, $\textbf{W}^{(l)}_{6}$ and $\textbf{W}^{(l)}_{7}$ are linear transformations. The output matrix of all entity embeddings can be represented as $\textbf{E}^{\mathcal{H}}_{t}$.
Table 2. Statistics of ICEWS14s, ICEWS18, ICEWS05-15 and GDELT datasets.
Dataset Entities $\mathcal{E}$ Relations $\mathcal{R}$ Training Facts Validation Facts Testing Facts Timestamps $\mathcal{T}$ Time Granularity
ICEWS14s 7,128 230 74,845 8,514 7,371 365 1 day
ICEWS18 23,033 256 373,018 45,995 49,545 304 1 day
ICEWS05-15 10,488 251 368,868 46,302 46,159 4,017 1 day
GDELT 7,691 240 1,734,399 238,765 305,241 2976 15 mins
3.5. Decoder
To achieve comprehensive predictions by leveraging entity and relation information from TKGs, $\mathsf{HisRes}$ adopts ConvTransE (Shang et al., 2019) as the score function, a widely used decoder in TKG reasoning, defined as follows:
(12) $p(o|s,r,t,\mathcal{G}_{0:t-1})=softmax(\mathbf{ConvTransE}(\textbf{s}^{\phi},\textbf{r}){\textbf{E}^{\phi}_{t}})\textnormal{,}$
(13) $\textbf{E}^{\phi}_{t}=\Theta^{\mathcal{H}}\textbf{E}^{\mathcal{H}}_{t}+(1-\Theta^{\mathcal{H}})\textbf{E}_{t}{,}$
(14) $\Theta^{\mathcal{H}}=\sigma(\textbf{W}_{8}\textbf{E}^{\mathcal{H}}_{t}+\textbf{b}_{gg})$
where $\textbf{W}_{8}$ and $\textbf{b}_{gg}$ denote linear layer and bias, $\textbf{s}^{\phi}\in\textbf{E}^{\phi}_{t}$, and $\textbf{E}^{\phi}_{t}$ represents the combined output of self-gating from the multi-granularity evolutionary encoder and global relevance encoder. It is worth noting that the relation prediction is calculated similarly.
3.6. Learning Objective
Entity prediction in KGs can be regarded as a multi-classification task. Hence, $\mathsf{HisRes}$ employs Cross-Entropy as the loss function. Importantly, we optimize the relation prediction to enhance relation representations, akin to TiRGN (Li et al., 2022c) and LogCL (Chen et al., 2024), to further enhance the performance of entity prediction. In addition, we introduce a coefficient $\alpha$ to control the weight of the two tasks. The learning objective is formalized as follows:
(15) $\mathcal{L}$ $=\alpha\sum_{(s,r,t)\in\mathcal{Q}^{e}_{t}}y^{e}_{t}\log p(o|s,r,t,\mathcal{G}_{0:t-1})$
$+(1-\alpha)\sum_{(s,o,t)\in\mathcal{Q}^{r}_{t}}y^{r}_{t}\log p(r|s,o,t,\mathcal{G}_{0:t-1})\textnormal{,}$
where $\mathcal{Q}^{e}_{t}$ and $\mathcal{Q}^{r}_{t}$ represent the queries of entities and relations at time $t$, and $y^{e}_{t}$ and $y^{r}_{t}$ denote the truth labels for the two tasks.
4. Experiments
In this section, we perform extensive experiments on four event-based TKG datasets to demonstrate the effectiveness and state-of-the-art performance of our model, which can be summarized in the following questions.
  • Q1: How does $\mathsf{HisRES}$ perform compared to state-of-the-art TKG reasoning studies?
  • Q2: How do the multi-granularity evolutionary encoder and global relevance encoder affect the performance of $\mathsf{HisRES}$?
  • Q3: How do the adaptive fusion of different entity representations and the updating of relations affect the performance of $\mathsf{HisRES}$?
  • Q4: How does the proposed superior ConvGAT affect the performance of $\mathsf{HisRES}$?
  • Q5: How does the performance fluctuation of $\mathsf{HisRES}$ with different settings?
Table 3. TKG entity extrapolation results on ICEWS14s, ICEWS18, ICEWS05-15 and GDELT. The time-filtered MRR, H@1, H@3, and H@10 metrics are multiplied by 100. The best results are boldfaced and the second SOTAs are underlined.
Model ICEWS14s ICEWS18 ICEWS05-15 GDELT
MRR H@1 H@3 H@10 MRR H@1 H@3 H@10 MRR H@1 H@3 H@10 MRR H@1 H@3 H@10
DistMult 15.44 10.91 17.24 23.92 11.51 7.03 12.87 20.86 17.95 13.12 20.71 29.32 8.68 5.58 9.96 17.13
ComplEx 32.54 23.43 36.13 50.73 22.94 15.19 27.05 42.11 32.63 24.01 37.50 52.81 16.96 11.25 19.52 32.35
ConvE 35.09 25.23 39.38 54.68 24.51 16.23 29.25 44.51 33.81 24.78 39.00 54.95 16.55 11.02 18.88 31.60
ConvTransE 33.80 25.40 38.54 53.99 22.11 13.94 26.44 42.28 33.03 24.15 38.07 54.32 16.20 10.85 18.38 30.86
RotatE 21.31 10.26 24.35 44.75 12.78 4.01 14.89 31.91 24.71 13.22 29.04 48.16 13.45 6.95 14.09 25.99
RE-NET 36.93 26.83 39.51 54.78 29.78 19.73 32.55 48.46 43.67 33.55 48.83 62.72 19.55 12.38 20.80 34.00
CyGNet 35.05 25.73 39.01 53.55 27.12 17.21 30.97 46.85 40.42 29.44 46.06 61.60 20.22 12.35 21.66 35.82
xERTE 40.02 32.06 44.63 56.17 29.31 21.03 33.51 46.48 46.62 37.84 52.31 63.92 19.45 11.92 20.84 34.18
RE-GCN 41.75 31.57 46.70 61.45 32.62 22.39 36.79 52.68 48.03 37.33 53.90 68.51 19.69 12.46 20.93 33.81
CEN 43.34 33.18 48.49 62.58 32.66 22.55 36.81 52.50 - - - - 21.16 13.43 22.71 36.38
TiRGN 44.61 33.90 50.20 64.89 33.66 23.19 37.99 54.22 50.04 39.25 56.13 70.71 21.67 13.63 23.27 37.60
CENET 39.02 29.62 43.23 57.49 27.85 18.15 31.63 46.98 41.95 32.17 46.93 60.43 20.23 12.69 21.70 34.92
RETIA 42.76 32.28 47.77 62.75 32.43 22.23 36.48 52.94 47.26 36.64 52.90 67.76 20.12 12.76 21.45 34.49
RPC - - - - 34.91 24.34 38.74 55.89 51.14 39.47 57.11 71.75 22.41 14.42 24.36 38.33
LogCL 48.87 37.76 54.71 70.26 35.67 24.53 40.32 57.74 57.04 46.07 63.72 77.87 23.75 14.64 25.60 42.33
$\mathsf{HisRES}$ 50.48 39.57 56.65 71.09 37.69 26.46 42.75 59.70 59.07 48.62 65.66 78.48 26.58 16.90 29.07 46.31
[…]
4.1. Experimental Setup
4.1.1. Datasets.
We conduct experiments on four real-world event-based TKGs extensively utilized in prior studies. Among them, three pertain to the Integrated Crisis Early Warning System (ICEWS) (Boschee et al., 2015): ICEWS14s, ICEWS18, ICEWS05-15. For comprehensive comparison and demonstrating the applicability and robustness of $\mathsf{HisRES}$, the Global Database of Events, Language, and Tone (GDELT) (Leetaru and Schrodt, 2013) dataset is also incorporated. Consistent with benchmarks (Li et al., 2021b; Chen et al., 2024), we split datasets into training (80%), validation (10%), and testing (10%) subsets based on their chronological sequences. Table 2 provides additional details on the statistical attributes of these datasets.
4.1.2. Baselines.
We compare our proposed $\mathsf{HisRES}$ with static KG reasoning methods (Yang et al., 2015; Trouillon et al., 2016; Dettmers et al., 2018; Shang et al., 2019; Sun et al., 2019) and the state-of-the-art TKG extrapolation models, which are shown below:
  • RE-NET (Jin et al., 2020) utilizes a combination of graph-based and sequential modeling to capture both temporal and structural interactions.
  • CyGNet (Zhu et al., 2021) constructs a historical vocabulary associated with all past events and rectifying predictions based on these events.
  • xERTE (Han et al., 2021) proposes a time-aware GAT that leverages temporal relations to explore subgraphs using temporal information.
  • RE-GCN (Li et al., 2021b) incorporates GNN and a recurrent unit to capture both the structural and temporal dependencies of TKGs, and introduces a static graph to augment initial embeddings.
  • CEN (Li et al., 2022a) employs a length-aware CNN within a curriculum learning and an online configuration to dynamically accommodate shifts in evolutionary patterns.
  • TiRGN (Li et al., 2022c) proposes time-aware decoders and incorporates a global historical vocabulary to predict based on whole history.
  • CENET (Xu et al., 2023) introduces contrastive learning into TKGs and designs a classifier to distinguish historical or global events.
  • RETIA (Liu et al., 2023) combines line graphs and captures the spatial-temporal patterns of entities and relations simultaneously.
  • RPC (Liang et al., 2023) incorporates a rule-based line graph for relations and mines graph structural correlations and periodic temporal interactions via two correspondence units.
  • LogCL (Chen et al., 2024) proposes an entity-aware attention encoder to capture crucial facts from local and global history.
4.1.3. Implementation Details.
The initial embedding dimension $d$ is configured to 200 for all datasets. The dropout rate of 0.2 is applied uniformly across all modules. The layers of GNNs are set to 2. Additionally, we utilize the Adam optimizer with a learning rate parameterized to 0.001. For the optimal historical length $l$ in multi-granularity evolutionary encoder, we conduct grid research according to the validation results, and the best value of $l$ is set to 9, 9, 10 and 7 for ICEWS14s, ICEWS18, ICEWS05-15 and GDELT, respectively. Noted that, to fair comparison, the task coefficient $\alpha$ is set to 0.7, and the static graph learning module is employed on ICEWS datasets similar to our baselines (Li et al., 2021b, 2022c; Liang et al., 2023; Chen et al., 2024). All baseline results are reported by previous works. We adopt a two-phase forward propagation strategy for the raw and inverse query sets similar to LogCL (Chen et al., 2024). All experiments are conducted on NVIDIA A800 GPUs.
4.1.4. Evaluation Metrics.
We report the widely-used time-aware filtered metrics (Han et al., 2021; Li et al., 2022c; Liang et al., 2023; Chen et al., 2024) (MRR and Hits@1,3,10) with percentage form in this paper.
4.2. Results (RQ1)
The performance of $\mathsf{HisRES}$ and advanced baselines under time-aware filtered metrics is reported in Table 3.
To comprehensively assess the performance of $\mathsf{HisRES}$, we initially compare it with static KG reasoning methods (i.e., DistMult, ComplEx, ConvE, ConvTransE, RotatE). As existing TKG baselines consistently outperform these methods across all benchmarks, underscoring the importance of considering temporal information and modeling evolutionary patterns within TKGs. $\mathsf{HisRES}$ achieves more progressive improvements with a 58.24% increase in MRR compared to ConvE. Furthermore, to provide valuable insights and compelling evidence addressing Q1, we compare $\mathsf{HisRES}$ with the best-performing TKG models. It demonstrates that $\mathsf{HisRES}$ is the state-of-the-art method, exhibiting a significant average improvement of 6.11%, 8.41%, 6.54% and 3.69% in time-filtered MRR, Hits@1, Hits@3 and Hits@10, respectively, across four event-based datasets.
Specifically, CyGNet and CENET focus on statistical analysis of historically repetitive facts but tend to overlook capturing structural interactions among snapshots. In contrast, $\mathsf{HisRES}$ not only models the structural and temporal dependencies of recent snapshots, but also incorporates a more comprehensive understanding of the entire history with the proposed global relevance encoder. It is more effective than simply rectifying predicted results.
In comparison to a series of approaches (Jin et al., 2020; Li et al., 2021b, 2022a, 2022c; Liu et al., 2023; Liang et al., 2023) that mainly focus on modeling each single snapshots from the most recent history, $\mathsf{HisRES}$ additionally considers the multi-granularity recent history. Concretely, we learn the interaction of each adjacent snapshot simultaneously (across time) while utilizing multi-hop links of GNNs to correlate the evolution of events intra-snapshot. Besides, we introduce a novel perspective that exclusively concentrates on relevant events throughout the entire history and further filters important links from them through ConvGAT. Although TiRGN considers global historical information, it still faces challenges similar to CyGNet, as it only employs a simple vector to represent global repetitive facts.
Furthermore, $\mathsf{HisRES}$ surpasses LogCL which achieves the second-best results and also provides a global perspective on top of recent snapshot modeling. This outstanding performance stems from multi-granularity and considering a self-gating mechanism to adaptively merge entity representations from different historical periods. Meanwhile, a more efficient approach to handling globally relevant facts is also helpful. In detail, we focus solely on facts directly relevant to the queries, thus reducing the number of historical facts. Additionally, integrating attention calculations into ConvGAT assists in selecting high-impact links or facts.
It should be noted that previous methods have not achieved significant improvements on GDELT due to the finer granularity of the time interval per snapshot (15 minutes), which makes GDELT more time-sensitive. $\mathsf{HisRES}$ provides an effective way of capturing multi-granularity features from both local evolution and global history while achieving a significant improvement of 11.92%, 15.44%, 13.55% and 9.40% in MRR, Hits@1, Hits@3, and Hits@10 on GDELT. This highlights the ability of $\mathsf{HisRES}$ to capture correlations within highly time-sensitive datasets, which have the typical characteristics of TKGs.
Table 4. Ablation studies with time-filtered metrics on ICEWS14s and ICEWS18.
Model ICEWS14s ICEWS18
MRR H@1 H@3 H@10 MRR H@1 H@3 H@10
$\mathsf{HisRES}$ 50.48 39.57 56.65 71.09 37.69 26.46 42.75 59.70
$\mathsf{HisRES}$-w/o-$\mathcal{G}$ 45.48 34.76 50.94 65.72 29.16 18.45 33.17 50.61
$\mathsf{HisRES}$-w/o-$\mathcal{G}^{\mathcal{H}}$ 41.83 31.49 47.01 61.74 31.55 21.53 35.41 51.48
$\mathsf{HisRES}$-w/o-MG 49.67 38.95 55.55 70.11 36.31 25.11 41.09 58.49
$\mathsf{HisRES}$-w/o-$\textit{SG}^{1}$ 50.04 39.34 55.86 70.28 37.08 25.76 42.07 59.39
$\mathsf{HisRES}$-w/o-$\textit{SG}^{2}$ 50.10 39.42 56.24 70.07 36.99 25.70 41.95 59.39
$\mathsf{HisRES}$-w/o-RU 50.17 39.37 56.17 70.38 36.99 25.79 41.79 59.12
$\mathsf{HisRES}$-w/-CompGCN 48.75 37.71 54.70 69.73 36.37 25.34 41.06 58.21
$\mathsf{HisRES}$-w/-RGAT 47.99 36.95 53.94 69.18 35.68 24.58 40.30 57.72
4.3. Ablation Studies (RQ2, RQ3 and RQ4)
To investigate the impact of different components within $\mathsf{HisRES}$ (Q2 and Q3), as well as the superiority of proposed ConvGAT (Q4), we conduct a series of ablation studies with ICEWS14s and ICEWS18 as representative datasets. The findings and results are summarized in Table 4.
  • The impact of different encoders (RQ2). To evaluate the indispensable local and global relevance structuring in TKG reasoning, we introduce two variants: $\mathsf{HisRES}$-w/o-$\mathcal{G}$ and $\mathsf{HisRES}$-w/o-$\mathcal{G}^{\mathcal{H}}$, indicating the removal of the multi-granularity evolutionary encoder and the global relevance encoder, respectively. The former highlights the importance of capturing evolutionary patterns from different granularities within the most recent snapshots. Conversely, the latter emphasizes the importance of additionally capturing global dependencies apart from local patterns, which notably enhances TKG reasoning.
  • The impact of self-gating and relation updating (RQ3). To investigate the effectiveness of adaptively merging different entity representations of TKGs, we replace self-gating with a simple summation. The variants are denoted as $\mathsf{HisRES}$-w/o-$\textit{SG}^{1}$ and $\mathsf{HisRES}$-w/o-$\textit{SG}^{2}$, corresponding to the two-stage fusion (equation (8) and (13)) in our methodology. The decreased performance shown in the second part of Table 4 indicates that the self-gating mechanism is effective for either type of entity representation fusion. Additionally, our strategy of relation updating (represented as $\mathsf{HisRES}$-w/o-RU) achieves an average improvement of 1.26% in MRR, underscoring the importance of considering the evolution of relations during aggregations.
  • To assess the superiority of the proposed ConvGAT, we replace it with two widely used GNN aggregators, CompGCN (Vashishth et al., 2020) and RGAT (Nathani et al., 2019). The resulting variants are denoted as $\mathsf{HisRES}$-w/-CompGCN and $\mathsf{HisRES}$-w/-RGAT, respectively, as shown in the third part of Table 4. Compared to the second-best result, the original $\mathsf{HisRES}$ achieves an average improvement of 3.59%, 4.68%, 3.84% and 2.26% in MRR, Hits@1, Hits@3 and Hits@10, respectively. It illustrates the significance of assigning different degrees of attention to various historical facts and underscores the efficiency of ConvGAT in calculating attention scores.
4.4. Sensitivity Analysis (RQ5)
To further explore the sensitivity of $\mathsf{HisRES}$ in different settings (Q5), we conduct experiments to explore the influence of (a) granularity level and (b) hidden layers of GNNs on the performance of $\mathsf{HisRES}$, which is shown in Figure 5.
Figure 5. Sensitivity analysis on ICEWS14s dataset.
4.4.1. Exploring the Influence of Granularity Level
The granularity level plays a pivotal role in controlling the interaction range inter-snapshot in the proposed multi-granularity evolutionary encoder simultaneously. It determines the quality of entity representations studied from the most recent snapshots. As illustrated in Figure 5(a), $\mathsf{HisRES}$ exhibits a compelling performance with capturing correlations at each 2 adjacent snapshots level. These findings highlight the impressive ability of $\mathsf{HisRES}$ to adapt to historical information from varying time spans. It is worth noting that $\mathsf{HisRES}$ maintains strong performance with different levels of granularity, demonstrating the robustness and effectiveness of our basic architecture.
4.4.2. Exploring the Influence of Hidden Layer of GNNs
Since $\mathsf{HisRES}$ integrates CompGCN and ConvGAT to learn multi-granularity evolutionary and globally relevant patterns, respectively, it is essential to evaluate the hidden layer numbers by investigating the influence of aggregation multi-hop neighbor information on entity representation. The results are depicted in Figure 5(b). Specifically, the two-hop results outperform both the one-hop and three-hop results. This observation suggests that overly simplistic neighborhoods or increasing the number of hops does not significantly enhance the performance of aggregating relevant events. Conversely, it may lead to oversmoothing or introduce unnecessary features that impede effective representations.
5. Conclusion
This paper introduces $\mathsf{HisRES}$, an innovative model for TKG reasoning. It comprises two novel encoders that excel at learning local and global entity representations by structuring historically relevant events. Specifically, we propose a multi-granularity evolutionary encoder to capture spatial-temporal interactions from inter-snapshot and intra-snapshot perspectives. In particular, we introduce a global relevance encoder featuring a distinctive ConvGAT to attend to crucial dependencies throughout the entire history. Furthermore, $\mathsf{HisRES}$ incorporates a self-gating mechanism for comprehensive fusion of TKG information. Experimental results on four event-based TKG datasets and various ablation studies demonstrate the effectiveness and superior performance of $\mathsf{HisRES}$. Future research avenues include exploring pruning techniques for global relevance and enhancing the modeling of temporal information.
Acknowledgements.